OpenStudy (anonymous):

Just need first step on factoring this 2x^2+16x+24 would you divide everything by 2?

OpenStudy (anonymous):

@jjuden @HeatDude @hartnn @michelle_DrNiah

OpenStudy (anonymous):

Yesh.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is this multiple choice?

OpenStudy (anonymous):

okay then what? lol i thought I needed the first step but Idk what to do fter that

OpenStudy (anonymous):

no

OpenStudy (anonymous):

When you divided it by two, you should have gtten 2(x^2+8x+12). YOu just factor that. What two factors of 12 will add up to 8? 3 and for? 12 and 1? Or 6, and 2?

OpenStudy (anonymous):

k, well you have to factor 2x^2+16x+24 2 is the only number here that can be factored. so 2(find what multiplied by 2 gives your original numbers) 2(x^2+8x+12) like @This_Is_Batman said

OpenStudy (anonymous):

k im good now thanx !!!!!!!!!!!!!!!!

OpenStudy (anonymous):

can you help me with one more thing?

OpenStudy (anonymous):

Bring it on. c:

OpenStudy (anonymous):

I need solutions to this > 1/2x^2+2x+3=0 can you please include steps and an explanation ????

OpenStudy (anonymous):

i need asap if you could please

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

Hang on, I'm trying.

OpenStudy (anonymous):

@whiterabbit ??

OpenStudy (anonymous):

I think you might have to do the opposite of b thing on this one.

OpenStudy (anonymous):

use the quadratic equation (or complete the square) since this polynomial can not be factored with real numbers

OpenStudy (anonymous):

That's what I thought. That's: \[\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\] Right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For your problem a would be 1/2, b would be 2, and c=3. Just plug in and solve. :)